Central High School has 480 freshmen. Of those freshmen, 333 take Algebra, 306 take Biology, and 188 take both Algebra and Biology. How many freshmen at Central High School take neither Algebra nor Geometry?
Found 2 solutions by jsmallt9, ewatrrr:
Answer by jsmallt9(3758)

The 333 who take Algebra includes both those who take Algebra and Biology and those who just take Algebra. With 188 who take both, this means that 333-188 or 145 freshmen take just Algebra.

The 306 who take Biology includes both those who take Algebra and Biology and those who just take Biology. With 188 who take both, this means that 306-188 or 118 freshmen take just Biology.

So we have 145 who just take Algebra, 118 who just take Biology and 188 who take both. This adds up to 451 freshmen. The other freshmen must take neither:
480-451 = 29 freshmen who take neither.

Answer by ewatrrr(24785)


Hi,
480 freshman
333 take Algebra, 306 take Biology, and 188 take both
333 + 306 - 188 = 451
480 - 451 = 29 took neither Algebra nor Geometry